Circular No. 8048

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003J, 2003K, 2003L, 2003M Discoveries of four apparent supernovae have been reported from un-

filtered CCD images: SN 2003J in NGC 4157 by R. Kushida (cf. IAUC 7890; via S. Nakano) and by T. Puckett and J. Newton (cf. IAUC 8043); SN 2003K in IC 1129 by Puckett and A. Langoussis (0.50-m reflector); SN 2003L in NGC 3506 by T. Boles (cf. IAUC 8030) and by M. Schwartz, B. Swift, and W. Li of LOTOSS (cf. IAUC 7906); and SN 2003M in UGC 7224 also by Swift and Li. The tabulated data below for SNe 2003J and 2003L are from Puckett and Boles, respectively.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003J Jan. 11.32 12^{h}10^{m}57.^{s}72 +50^{o}28′31′′.8 16.7 65′′.5 W, 32′′.1 S
2003K Jan. 11.32 15 32 01.91 +68 14 35.9 15.4 6′′.1 E, 10′′.1 S
2003L Jan. 12.15 11 03 12.33 +11 04 38.3 16.9 9′′.0 W, 1′′.5 N
2003M Jan. 13.5 12 13 21.00 +21 38 48.4 17.3: 38′′.2 E, 44′′.1 N

Additional magnitude estimates, unfiltered CCD data except for the Sky Survey [B = Boles; K = Kushida; L = KAIT/LOTOSS; O = T. Kobayashi, Oizumi, Gunma, Japan, 0.41-m reflector (also via Nakano); P = Puckett; S = Palomar Sky Survey; and T = Tenagra II/LOTOSS]: SN 2003J, 1995 Apr. 4, [21.0: (S, P); 2001 June 24, [20.5: (P); 2002 May 25, [18: (K); 2003 Jan. 11.857 UT, 16.3 (K); 12.66, 16.3 (K). SN 2003K, 1991 May 17, [21.0: (S, P); 2002 Aug. 4, [20.2: (P). SN 2003L, 2002 Apr. 4, [19.0 (B); Nov. 4, [18.5 (B); Nov. 14.5, [19.5: (T); 2003 Jan. 6.700, 17.3 (O); 13.49, 16.0: (T); 13.55, 16.0: (L). SN 2003M, 2002 May 14.3, [19.0: (L); Dec. 23.5, 18.5: (L). Position end figures for SN 2003J from Kushida’s image (measured by Y. Koshida): 57^{s}.52, 32′′.7. Puckett adds that both SN 2003J and SN 2003K were confirmed on CCD frames taken on Jan. 12.15. LOTOSS position end figures for SN 2003L: 12^{s}.32, 38′′.0.

K. Ayani, Bisei Astronomical Observatory (BAO); and T. Hashimoto and H. Yamaoka, Kyushu University, write that preliminary reduction of a low-resolution spectrum (range 400–800 nm) of SN 2003J, obtained with the BAO 1.01-m telescope on Jan. 12.8 UT, shows prominent Hα emission (FWHM ∼ 10000 km/s) with a P-Cyg profile, suggesting that it is a normal type-II supernova. Ayani and Yamaoka add that another low-resolution spectrum (obtained as above) of SN 2003K indicates that it is a type-Ia supernova several days before maximum light; adopting the NASA/IPAC Extragalactic Database recession velocity of 6540 km/s, the expansion velocity deduced from Si II (rest 635.5 nm) is ∼ 12100 km/s.

2003 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green